OpenStudy (anonymous):

A disc jokey has 10 songs to play. Four are slow songs, and six are fast songs. Each song is to be played only once. In how many ways can the disc jokey play the 10 songs if :The songs can be played in any order :The first song must be a slow song and the last song must be a slow song. :The first two songs must be fast songs.

ganeshie8 (ganeshie8):

10 songs in a sequence can be played in 10! ways

ganeshie8 (ganeshie8):

For the second q- 1) you fix first and last as slow songs : 4 * 3 2) rest of the 8 songs can be arranged in 8! ways so total ways he can play by having first & last as slow songs = 4*3*8!

ganeshie8 (ganeshie8):

does that help ? last q can u try.. .

OpenStudy (anonymous):

i donno thats kinda confusing to me...

ganeshie8 (ganeshie8):

did u get my method... how i worked out first two ?

OpenStudy (anonymous):

can u explain it a bit more how u did that?

ganeshie8 (ganeshie8):

10 different songs, so think 10 spaces to fill with songs ok _ _ _ _ _ _ _ _ _ _

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

now lets work out how many different ways you can play these 10 songs

ganeshie8 (ganeshie8):

First song can be played by choosing any of the 10 songs. Second song can be played by choosing any of the remaining 9 songs. (its not 10 becos, repitition is not allowed, so we cant play the same song. one less)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

similarly, Third song can be played in 8 ways Fourth song can be played in 7 ways ...... Ninth song can be played in 2 ways Tenth song can be played in 1 ways total ways = 10 * 9 * 8 * 7 * ...... 2*1 = 10!

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

above is the fundamental thing u should understand. if u understand this, you can do ANY permutation problem!

ganeshie8 (ganeshie8):

did u get the method ? second q- is little different but you can use the same method there also

OpenStudy (anonymous):

i think so

ganeshie8 (ganeshie8):

great ! 2nd q : :The first song must be a slow song and the last song must be a slow song.

ganeshie8 (ganeshie8):

if you forget about first and last spaces for a while. middle 8 spaces can be arranged in how many ways ?

ganeshie8 (ganeshie8):

8! right ?

OpenStudy (anonymous):

o ya duh

ganeshie8 (ganeshie8):

first and last you fixing with two slow songs. so remaining 8 middle spaces can be arranged in 8! ways. its same logic we used in the first q S _ _ _ _ _ _ _ _ S

ganeshie8 (ganeshie8):

we are not done yet. we have 4 slow songs. so first & last space can be chosen in 4*3 ways so total ways = 4*3*8!

ganeshie8 (ganeshie8):

makes sense ? let me know if it doesnt... third q is same as 2nd one

OpenStudy (anonymous):

o ya ok

ganeshie8 (ganeshie8):

:The first two songs must be fast songs.

ganeshie8 (ganeshie8):

F F _ _ _ _ _ _ _ _

ganeshie8 (ganeshie8):

remaining 8 spaces can be filled in how many ways can u tell me

OpenStudy (anonymous):

8 are slow

ganeshie8 (ganeshie8):

hmm read the question once. there are only 4 slow songs :P

ganeshie8 (ganeshie8):

you simply fix first two as Fast. remaining 8 spaces can be filled in how many ways ?

OpenStudy (anonymous):

3 ways

ganeshie8 (ganeshie8):

nope. 8 spaces & 8 songs. so its 8! ways right ?

OpenStudy (anonymous):

8 ways

OpenStudy (anonymous):

yes i already said that

ganeshie8 (ganeshie8):

since there are 6 Fast songs, first two Fast songs can be selected from them in : 6*5 ways. so total ways = 6*5*8!

ganeshie8 (ganeshie8):

its not 8 ways. its 8! ways = 8*7*6*... 2*1

ganeshie8 (ganeshie8):

8 \(\ne \) 8!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

why do you have exclamation marks after 8?

ganeshie8 (ganeshie8):

! is a symbol for factorial 4! = 4*3*3*1 3! = 3*2*1 8! = 8*7*6*5*4*3*2*1

OpenStudy (anonymous):

o ok

ganeshie8 (ganeshie8):

100! = 100*99*98...... 3*2*1 it just says, to perform series of multiplications as above

ganeshie8 (ganeshie8):

for the 3rd q- we got answer as : 6*5*8! can you do the factorial and and see what value u get.. ?

OpenStudy (anonymous):

6= 6*5*4*3*2*1 5 = 5*4*3*2*1 8 = 8*7*6*5*4*3*2*1 so this is what u were doing and to get a vlue?

ganeshie8 (ganeshie8):

hmm no

ganeshie8 (ganeshie8):

notice that exclamation sigh is there only for 8

ganeshie8 (ganeshie8):

6*5*8! = 30*8! = 30 * 8*7*6*5*4*3*2*1

OpenStudy (anonymous):

o 5 * 6 = 30 * 8! = 30 * 8 * 7* 6*5*4*3*2*1 ok